What is the equation of the line that passes though points (4\5 , 1\5) and (1\2, 3\2)?
Answer by Cromlix(4381)

To find gradient:
y2 - y1/x2 - x1
3/2 - 1/5/1/2 - 4/5
13/10 / -3/10
=130/-30
=13/-3
Using equation:
y - b = m(x - a)
a = 1/2, b = 3/2 and m = -13/3
y - 3/2 = -13/3(x - 1/2)
y = -13/3x + 22/6
Or
Multiply thro' by 6
6y = -26x + 22
Hope this helps
:-)